OpenStudy (anonymous):

(-7x(squared)+ 5x-4) - (6x(squared)-8)

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no problem

OpenStudy (turingtest):

rather that write the word "squared" every time it is better to use the ^ notation: (-7x^2+5x-4)-(6x^2-8)

OpenStudy (anonymous):

oh okay

OpenStudy (turingtest):

first thing is to distribute the negative sign to get rid of the parentheses

OpenStudy (anonymous):

yeah

OpenStudy (turingtest):

remember that\[-(ax^2-bx+c)=-1(ax^2-bx+c)=-ax^2+bx-c\]so do the same thing to the second set of parentheses in your expression what do you get?

OpenStudy (anonymous):

I tried doing that before and I got a really random answer